OpenStudy (anonymous):

Determine the coefficient of x^4y^5 in the expansion of (3x - 2y)^9.

OpenStudy (anonymous):

is the answer -326592

OpenStudy (phi):

Yes, it is \[\left(\begin{matrix}9 \\ 5\end{matrix}\right)\cdot 3^4\cdot(-2)^5\]